Represent the following situations in the form of quadratic equations: 
The product of two consecutive positive integers is 306. We need to find the integers. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Let us consider,
The first integer number = x
Thus, the next consecutive positive integer will be = x + 1
Product of two consecutive integers = x × (x +1) = 306
⇒ x^2+ x = 306
⇒ x^2+ x – 306 = 0
Therefore, the two integers x and x+1, satisfies the quadratic equation, x^2+ x – 306 = 0

Thanks